Az első valószínűségszámítás dolgozat eredménye

1.  2.  3.  4.  5.

diákok:

Bordi István  ±  −  +  −  −±
Boros László  −  +  +  +  +
Böszörményi Zsolt  −  −  +  +  −
Csipkés László  −  −  +  +  ±
Ferenczik Zoltán  +  +  +  +  +
Fogas János  −  −  +  +  −
Havasi Balázs  −  −  +  −  −
Hennel Zoltán  −  −  +  −  −
Illés József  +  +  +  ±  ±
Jágó Jácint  −  +  +  +  −
Kiss Róbert  −  +  +  +  +±
Lászlók Zsolt  −  ±  +  +  −
Lázár Gábor  −  −  +  +  +
Mezei Tamás  +  +  +  +  ±
Molnár Adrienn  −  +  +  +  −
Molnár Zoltán  −  +  +  −  ±
Nagy András  +  +  +  −  +
Paper Andrea  −  +  +  +  −
Pelejtei Gergely  −  −  +  +  ±
Szalontai Gábor  +  +  +  +  ±
Tordai Jácint  −  +  +  +  +
Ujjobbágy Attila  −  −  −  −  −
Vágner József  −  −  +  +  ±
Virág Gabriella  +  +  +  +  +
Zádor István  −  −  +  −  −